# Date of publication xxxx 00, $0$, date of current version xxxx 00, $0$. n/a

### v1

@@ Abstract
The shift from a linear to a circular economy has the potential to simultaneously reduce uncertainties of material supplies and waste generation. To date, the development of robotic and, more generally, autonomous systems have been rarely integrated into circular economy implementation strategies. In this review, we merge deep-learning vision, compartmental dynamical thermodynamics, and robotic manipulation into a theoretically-coherent physics-based research framework to lay the foundations of circular flow designs of materials, and hence, to speed-up the transition from linearity to circularity. Then, we discuss opportunities for robotics in circular economy.
@@ I Introduction
Visualize for a moment that a built environment and human activities are contained within a plenipotentiary unit or compartment. This compartment has an input flow and an output flow of materials: the former comes from natural reserves, whereas the latter is the generation of waste and pollution. The central issue with the input flow of the compartment is that the supplies of several raw materials have a high level of risk and are also needed to manufacture high-tech and green technologies essential to a modern society [1]. Hence, these materials are labeled as “critical” by the European Union [2] and the United States [3].
@@
The circular economy (CE) paradigm aims in minimizing both the input and the output flows of the plenipotentiary unit or compartment through an increase of maintenance, reuse, remanufacturing, recycling, and material efficiency [8].
Our contributions: To increase the implementation of circularity strategies at both local and global scales, in this review:
  • We merge deep-learning vision, compartmental dynamical thermodynamics, and robotic manipulation into a theoretically-coherent research framework. Our framework has a double benefit: namely, it integrates robotics into the systemic perspective needed to address circular economic challenges; and it improves the scientific foundation of the theory of CE.
  • We discuss opportunities for robotics in a circular economy to reduce the existing gap between Industry 4.0 techniques and Sustainability [9].
The rest of the paper is organized as follows. Section II covers the literature review relevant for our theoretically-coherent research framework by focusing on four different topics. Then, in Section III, we present a unification of the four key topics discussed in Section II, whereas in Section IV we collect existing models from the literature aligned with our framework. In Section V, we focus on the opportunities for robotics in a circular economy. Finally, in Section VI, we give some conclusions.
@@ II THREE REPRESENTATIVE TOPICS FOR CIRCULAR ECONOMY
This section covers the literature regarding the four topics mentioned in the title and intersected in this review; namely, circular economy (CE), compartmental dynamical thermodynamics (CDT), robotic manipulation (RM), and deep-learning-vision (DLV).
@@ II-A Circular Economy
Velenturf et al. [10] view circular economy as “[…] an emerging practical ideology that lacks an evidence-based theoretical framework to guide implementation […]”. Kirchherr et al. [11] identified 114 definitions of circular economy to highlight the lack of clarity and agreement about its meaning. This lack of evidence-based and clear foundations is one of the main motivations for this review. Here, we adopt the popular and intuitive definition of the Ellen MacArthur Foundation [8], which looks at closing the loops of material flows similarly to what we discussed in Section I. The large number of definitions of CE goes in parallel with the large number of indicators developed to measure circularity [12, 13, 14]. For example, De Pascale et al. [15] identified 61 indicators.
@@ II-C Robotic Manipulation for Reprocessing
Besides grasping and manipulating tasks, often in automated systems in the circular economy contexts involve more articulated operations. One example is represented by disassembling elements [54], where complex bimanual operations are required; for example unscrewing, removing components, cutting, etc. [55, 56, 57]. Disassembly operations are becoming particularly important in waste electronic and electrical equipment (WEEE) [58]. Autonomous systems have been proposed for disassembly tasks of specific items, e.g., electronic components [59, 60], batteries [61], cellphones [62], press fitted components [63], etc. For disassembling tasks, multifunctional grippers enabling bimanual tasks have been presented in [64, 65].
@@ II-D Deep-Learning Vision for Robotics
The breakthroughs in deep learning over the last decade have significantly improved the accuracy, speed, and generalization power of vision systems. More recently, the robotics community is leveraging such advances to enhance the ability of robots to navigate and interact with environments more complex than research testbed laboratory environments. Lee et al. [70] addressed contact-rich tasks by training a deep neural network on RGB, depth, force-torque, and proprioception data. Vision-based navigation were developed in [71, 72]; the former used a deep-reinforcement-learning approach, whereas the latter combined model-based control with learning from images.
@@
Object tracking is a useful ability for a robot that needs to interact with a dynamic environment. Object tracking systems were developed in [73, 74, 75, 76, 77, 78]. A literature review on deep-learning-based visual tracking can be found in [79]. An effective technique for visual tracking is to use object detection on each video frame, but annotating data manually is very time consuming. Hence, Horváth et al. [80] automated the annotation process. Another important task for intelligent robots is estimating the pose of target objects, especially for their manipulation. Pose estimation of objects was developed in [81, 82, 83, 84].
@@
Visual servoing typically consists of a combination of one or more cameras, the image processing module, and the control algorithms that translate the image information and other sensor data into robot actions. Deep-learning visual servoing was proposed in [85, 86, 87]. Costante et al. [88] proposed a data-driven odometry module that provides information about the uncertainty of its estimations.
@@
Deep-learning vision has also been proposed for autonomous marine debris detection [89, 90, 91], whereas non-marine waste was the focus of [92, 93, 94, 95]. Ku et al. [37] used a deep neural network to find an effective grasping pose for sorting construction and demolition waste. Grasping waste items is particularly challenging since their shape is difficult to predict. In contrast, manufacturing processes have less uncertainties since the geometry of the products are known. A similar consideration holds for disassembly; the conditions of an end-of-life device are difficult to predict, which makes the automation of their disassembly sequence more challenging than assembling processes in manufacturing, where the assembly sequence is known.
@@
Jahanian et al. [96] focused on electronic waste for disassembly operations, whereas [97, 98, 99] detected the screws since their removal is the initial step of many disassembly sequences. Medical waste was considered in [100, 101]. Visual material recognition can improve both the effectiveness of robotic grasping (e.g., lower control gains with fragile items) and the automation of material sorting for a circular economy [102]. Material recognition from images was performed in [103, 104, 105].
@@
This representation of a life-cycle is referred to as “compartmental diagram” in [21]. The origin $O$ of the frame of reference is the tail of the position vector of an infinitesimal cube of material which crosses the compartments. The compartments are connected by material flows (the arrows) and the material type is indicated by $\beta_{q}$, with $q\in[1,Q]$. The total number of compartments is $n_{c}$. The “Transportation” compartment usually does not modify the material, hence the input flow is of the same type as the output flow. The recovery chain can be automated to reduce the costs of waste sorting and re-manufacturing operations [101]. The robot can be equipped with a camera and a gripper for handling waste materials or end-of-first-use products. Then, the reprocessed materials or products can reenter the supply chain at different stages; recycled material could be used by a manufacturer to reduce its dependence on finite natural resources, while a repaired product could be transported directly to a new consumer.
@@
Topic 2 (CDT): Along with circularity, the Rankine cycle has another important feature: its modeling begins by considering, in turn, the thermodynamics of each component (i.e., the turbine, condenser, pump, and boiler). Hence, the modeling is based on compartmental dynamical thermodynamics, which conceives, as discussed in Section II-B, multi-machine systems as a set of thermodynamic compartments connected by material flows [26]. Recall that the simplest analysis of a Rankine cycle is at steady state, i.e., when the rate of change of energy $\dot{E}=0$. Under the simplifying, but practically useful, assumption of steady state, the term “dynamical” in “compartmental dynamical thermodynamics” is removed since, at equilibrium, the dynamics of the system is not considered.
@@
Topic 4 (DLV): Consider again a manipulator as a thermodynamic compartment. If it is equipped with one or more cameras as in Fig. 1, then these deep-learning-vision algorithms can process the images in real-time to improve the ability of the robot-compartment to handle waste materials or end-of-first-use products.
@@ V OPPORTUNITIES FOR ROBOTICS IN CIRCULAR ECONOMY
Opportunities in application of robot technologies in a circular economy are wide-ranging satisfy two key aspects connected with waste management: $i$) substitution of human activities in the case of elevated health risk due to the nature of the process or to the waste material properties; and $ii$) performance increment in waste sorting outputs or waste collection quality, avoiding high costs due to destination of improper material to the wrong waste treatment process in the new prospective of a circular economy. Waste separated collection [122] and waste sorting were widely investigated with automation technologies focusing on the high level of waste flows to be managed and optimizing the high margins with respect to domestic waste management. The promising prospects of reducing management costs and increasing the effectiveness and efficiency of selection processes has stimulated the in-depth study of innovative robotization technologies [123], in the face of a notable quantity of industrial automation solutions on a market scale with sustainable solutions applied to municipal solid waste [124, 125].
The transition from linear waste management to a circular economy requires a notable leap in quality in the performance of all phases of the process, in order to avoid effects of economic inefficiency due to the multiplication of transport and management costs due to flows affected by unsuitable materials with respect to the quality conditions required by recycling processes. From this perspective, robotization technologies can play a decisive role in completion of automatic management of supply chains, in the final classification of flows destined for recycling in the face of reduced levels of unwanted or contaminated materials, replacing processes that are still largely based on manual selection [126].
@@
VI CONCLUSION
In this paper, we coherently combined the literature related to four key topics: deep-learning vision, compartmental dynamical thermodynamics, robotic manipulation, and circular economy. Specifically, the latter is the area of application wherein the other three topics are presented.

### v2

@@ Abstract
The shift from a linear to a circular economy has the potential to simultaneously reduce uncertainties of material supplies and waste generation. However, to date, the development of robotic and, more generally, autonomous systems have been rarely integrated into circular economy implementation strategies despite their potential to reduce the operational costs and the contamination risks from handling waste. In addition, the science of circularity still lacks the physical foundations needed to improve the accuracy and the repeatability of the models. Hence, in this paper, we merge deep-learning vision, compartmental dynamical thermodynamics, and robotic manipulation into a theoretically-coherent physics-based research framework to lay the foundations of circular flow designs of materials. The proposed framework tackles circularity by generalizing the design approach of the Rankine cycle enhanced with dynamical systems theory. This differs from state-of-the-art approaches to circular economy, which are mainly based on data analysis, e.g., material flow analysis (MFA). We begin by reviewing the literature of the three abovementioned research areas, then we introduce the proposed unified framework and we report the initial application of the framework to plastics systems along with initial simulation results of reinforcement-learning control of robotic waste sorting. This shows the framework applicability, generality, scalability, and the similarity and difference between the optimization of artificial neural systems and the proposed compartmental networks. Finally, we discuss the still not fully exploited opportunities for robotics in circular economy and the future challenges in the theory and practice of the proposed circularity framework.
@@ I Introduction
Visualize for a moment that a built environment and human activities are contained within a unit or compartment. This compartment has an input flow and an output flow of materials: the former comes from natural reserves, whereas the latter is the generation of waste and pollution. The central issue with the input flow of the compartment is that the supplies of several raw materials have a high level of risk and are also needed to manufacture high-tech and green technologies essential to a modern society [1]. Hence, these materials are labeled as “critical” by the European Union [2] and the United States [3].
@@
In this context, the circular economy (CE) paradigm aims in minimizing both the input and the output flows of the unit or compartment through an increase of maintenance, reuse, remanufacturing, recycling, and material efficiency [8, 9]. Keeping materials in use within the economy (i.e., the unit) for a longer time reduces the extraction of finite resources and it also could increase the self-sufficiency of the communities [10]. According to a 2021 report from the Centre for European Policy Studies (CEPS), there are several barriers to the adoption of circularity practices such as existing policies, economic factors, supply chains, technology, and consumer preferences [11]. From a scientific and educational perspective, the principles of circularity are very rarely integrated into highly technical programs and courses, which further delays their adoption by the general public and younger generations [12, 13, 14].
One of the main reasons for the weak intersection between technical subjects and circularity is the lack of solid physical and engineering foundations of the latter since, traditionally, CE is a topic of business and environmental sciences. Hence, there is a societal need for integrating CE principles into technically rigorous programs while still maintaining their high mathematical and physics standards. Indeed, such standards and skills are essential to create accurate circularity models and to design and prototype circularity-oriented devices and intelligent systems.
In this context, this paper makes the following contributions:
  • We merge deep-learning vision, compartmental dynamical thermodynamics, and robotic manipulation into a theoretically-coherent research framework. Our framework has a double benefit: namely, it integrates robotics into the systemic perspective needed to address circular economic challenges; and it improves the physical foundations of the theory of CE.
  • We report the initial application of the proposed framework to plastics systems (Section VI).
  • We discuss opportunities for robotics in a circular economy to address the existing gap between Industry 4.0 techniques and Sustainability [15] and, from an educational point of view, to address the gap between robotics and environmental programs (Section V).
The rest of the paper is organized as follows. Section II covers the literature review relevant for our theoretically-coherent research framework by focusing on four different topics. Then, in Section III, we present a unification of the four key topics discussed in Section II, whereas in Section IV we collect existing models from the literature aligned with our framework. In Section V, we focus on the opportunities for robotics in a circular economy, while in Section VI we apply the proposed framework to plastics systems. Finally, in Section VII, we give conclusions.
@@ II THREE REPRESENTATIVE TOPICS FOR CIRCULAR ECONOMY
This section covers the literature regarding the four topics mentioned in the title and intersected in this paper; namely, circular economy (CE), compartmental dynamical thermodynamics (CDT), robotic manipulation (RM), and deep-learning vision (DLV).
@@ II-A Circular Economy
Velenturf et al. [16] view circular economy as “[…] an emerging practical ideology that lacks an evidence-based theoretical framework to guide implementation […]”. Kirchherr et al. [17] identified 114 definitions of circular economy to highlight the lack of clarity and agreement about its meaning. This lack of evidence-based and clear foundations is one of the main motivations for this paper. Here, we adopt the popular and intuitive definition of the Ellen MacArthur Foundation [8], which looks at closing the loops of material flows similarly to what we discussed in Section I. The large number of definitions of CE goes in parallel with the large number of indicators developed to measure circularity [18, 19, 20]. For example, De Pascale et al. [21] identified 61 indicators.
@@
To date, material flow analysis (MFA) is the main methodology for assessing circularity [27, 28, 29, 30]. It is essentially based on the analysis of real data of material stocks and flows recorded over time. In terms of its physical foundations, MFA is based on the mass balance and the MFA results are usually represented with Sankey diagrams with arrows indicating the direction of the flows between different life-cycle stages: e.g., extraction and manufacturing [27, 28, 29, 30]. With respect to MFA, the framework proposed in this paper, firstly, adds thermodynamic laws to the mass balances and applies them to the control surfaces used in the design of thermodynamic cycles; and secondly, it makes use of dynamical system models (i.e, ordinary differential and difference equations) instead of data modeling and analysis.
@@ II-C Robotic Manipulation for Reprocessing
Besides grasping and manipulating tasks, often in automated systems in the circular economy contexts involve more articulated operations. One example is represented by disassembling elements [64], where complex bimanual operations are required; for example unscrewing, removing components, cutting, etc. [65, 66, 67]. Disassembly operations are becoming particularly important in waste electronic and electrical equipment (WEEE) [68]. Autonomous systems have been proposed for disassembly tasks of specific items, e.g., electronic components [69, 70], batteries [71], cellphones [72], and press fitted components [73]. For disassembling tasks, multifunctional grippers enabling bimanual tasks have been presented in [74, 75].
@@ II-D Deep-Learning Vision for Robotics
The breakthroughs in deep learning over the last decade have significantly improved the accuracy, speed, and generalization power of vision systems. More recently, the robotics community is leveraging such advances to enhance the ability of robots to navigate and interact with environments more complex than research testbed laboratory environments. Lee et al. [80] addressed contact-rich tasks by training a deep neural network on RGB, depth, force-torque, and proprioception data. The real robot is controlled by a reinforcement learning (RL) policy learned through self-supervision to perform peg insertion. Vision-based navigation was developed in [81, 82]; the former used a deep RL approach, whereas the latter combined model-based control with learning from images.
@@
Object tracking is a useful ability for a robot that needs to interact with a dynamic environment. An object tracking system was developed by Huang et al. [83] using three novel modules, namely, the motion disentanglement network, the patch association network, and the patch memory network, while the system of Gordon et al. [84] simultaneously tracks and updates a deep object tracker in real-time with a single forward pass. Other algorithms were proposed in [85, 86, 87, 88]. A literature review on deep-learning-based visual tracking can be found in [89]. An effective technique for visual tracking is to use object detection on each video frame, but annotating data manually is very time consuming. Hence, Horváth et al. [90] automated the annotation process. Another important task for intelligent robots is estimating the pose of target objects, especially for their manipulation. Pose estimation of objects was developed in [91, 92, 93, 94].
@@
Visual servoing typically consists of a combination of one or more cameras, the image processing module, and the control algorithms that translate the image information and other sensor data into robot actions. Al-Shanoon et al. [95], for example, trained a convolutional neural network using only synthetic data and they used a feedback control scheme based on the 3-D object pose for manipulation tasks. Other approaches were proposed in [96, 97], while Costante et al. [98] proposed a data-driven odometry module that provides information about the uncertainty of its estimations.
Deep-learning vision has also been proposed for autonomous marine debris detection [99, 100, 101], whereas non-marine waste was the focus of [102, 103, 104, 105]. In particular, Bashkirova et al. [102] collected and published the first in-the-wild industrial-grade waste detection and segmentation dataset, namely, ZeroWaste, to facilitate the development of vision systems for highly cluttered recycling scenarios, while Kiyokawa et al. [105] opted for graspless push-and-drop and pick-and-release manipulation for their robotic waste sorter.
@@
Ku et al. [47] used a deep neural network to find an effective grasping pose for sorting construction and demolition waste. Grasping waste items is particularly challenging since their shape is difficult to predict. In contrast, manufacturing processes have less uncertainties since the geometry of the products are known. A similar consideration holds for disassembly; the conditions of an end-of-life device are difficult to predict, which makes the automation of their disassembly sequence more challenging than assembling processes in manufacturing, where the assembly sequence is known.
@@
Jahanian et al. [106] focused on electronic waste for disassembly operations, whereas [107, 108, 109] detected the screws since their removal is the initial step of many disassembly sequences. Medical waste was considered in [110, 111]. Visual material recognition can improve both the effectiveness of robotic grasping (e.g., lower control gains with fragile items) and the automation of material sorting and mapping for a circular economy as extensively reviewed and discussed by Zocco et al. [112] in proposing material measurement units, which are distributed autonomous visual recognizers of the materials located in a certain area of interest, e.g., a town. Material recognition from images was performed by Bell et al. [113]. Specifically, they introduced a large-scale open dataset of materials in the wild, namely, MINC, and then, they trained convolutional neural networks to simultaneously perform material recognition and segmentation. Other works on material recognition are [114, 115].
@@
This representation of a life-cycle is referred to as “compartmental diagram” in [31]. The origin $O$ of the frame of reference is the tail of the position vector of an infinitesimal cube of material which crosses the compartments. The compartments are connected by material flows (the arrows) and the material type is indicated by $\beta_{q}$, with $q\in[1,Q]$. The total number of compartments is $n_{c}$. The “Transportation” compartment usually does not modify the material, hence the input flow is of the same type as the output flow. The recovery chain can be automated to reduce the costs of waste sorting and re-manufacturing operations [111]. The robot can be equipped with a camera and a gripper for handling waste materials or end-of-first-use products. Then, the reprocessed materials or products can reenter the supply chain at different stages; recycled material can be used by a manufacturer to reduce its dependence on finite natural resources, while a repaired product can be transported directly to a new consumer.
@@
Topic 2 (CDT): Along with circularity, the Rankine cycle has another important feature: its modeling begins by considering, in turn, the thermodynamics of each component (i.e., the turbine, condenser, pump, and boiler). Hence, the modeling is based on compartmental dynamical thermodynamics, which conceives, as discussed in Section II-B, multi-machine systems as a set of thermodynamic compartments connected by material flows [36]. Recall that the simplest analysis of a Rankine cycle is at steady state, i.e., when the time rate of change of energy $\dot{E}=0$. Under the simplifying, but practically useful, assumption of steady state, the term “dynamical” in “compartmental dynamical thermodynamics” is removed since, at equilibrium, the dynamics of the system is not considered.
@@
Topic 4 (DLV): Consider again a manipulator as a thermodynamic compartment. If it is equipped with one or more cameras as in Fig. 1, then these deep-learning vision algorithms can process the images in real-time to improve the ability of the robot-compartment to handle waste materials or end-of-first-use products.
In summary, Fig. 1 merges the four topics as follows. The rectangles are the control surfaces of the compartments (this is CDT); the robot depicted at the top in the recovery chain (green arrows) is a compartment that could sort or repair waste items using advanced manipulation and machine vision (this includes RM and DLV in the diagram). Finally, CE is included in the diagram since the material moves from a compartment to another in a circular fashion in order to reduce the two flows indicated with the red arrows. The literature review of each topic was covered in Section II in the following order: CE, CDT, RM, and DLV. The next section will review models and systems proposed in the literature that derive from mass balances and/or thermodynamic laws, and hence, they can be considered compartments $c^{k}_{i,j}$ to be connected as in Fig. 1 to simulate existing and hypothetical circular flows.
@@ IV A BIBLIOTHECA OF COMPARTMENTS
This section has reviewed existing models in the literature that fit within the proposed research framework. The next section will focus specifically on robots, which are a particular class of compartments that could play a significant role in a circular economy as depicted in Fig. 1. The derivation of the standard form of robot dynamics from the first law of thermodynamics was demonstrated in [111], Proposition 1.
@@ V OPPORTUNITIES FOR ROBOTICS IN CIRCULAR ECONOMY
Opportunities in application of robot technologies in a circular economy are wide-ranging and satisfy two key aspects connected with waste management: $i$) substitution of human activities in the case of elevated health risk due to the nature of the process or to the waste material properties; and $ii$) performance increment in waste sorting outputs or waste collection quality, avoiding high costs due to destination of improper material to the wrong waste treatment process in the new prospective of a circular economy. Waste separated collection [132] and waste sorting were widely investigated with automation technologies focusing on the high level of waste flows to be managed and optimizing the high margins with respect to domestic waste management. The promising prospects of reducing management costs and increasing the effectiveness and efficiency of selection processes has stimulated the in-depth study of innovative robotization technologies [133], in the face of a notable quantity of industrial automation solutions on a market scale with sustainable solutions applied to municipal solid waste [134, 135].
The transition from linear waste management to a circular economy requires a notable leap in quality in the performance of all phases of the process, in order to avoid effects of economic inefficiency due to the multiplication of transport and management costs due to flows affected by unsuitable materials with respect to the quality conditions required by recycling processes. From this perspective, robotization technologies can play a decisive role in the completion of automatic management of supply chains, in the final classification of flows destined for recycling in the face of reduced levels of unwanted or contaminated materials, and replacing processes that are still largely based on manual selection [136].
@@
The next section will illustrate the application of the proposed framework to a case study of plastics including a robotic waste sorting compartment. This aligns with the current section, which identified waste sorting as a task to potentially automate with robots.
VI Framework Application to a Case Study and Challenges
VI-A Initial Application to Plastics Systems
This section covers the initial application of the proposed framework to design circular flows of plastics. Targeting plastics flows is motivated by the rising accumulation of plastic waste in natural environments [125]. This study considers also bio-plastics as they are made from renewable and carbon-sink feedstock, and hence, their use makes the communities more sustainable and resilient than with their synthetic counterpart [138]. A project recently funded by the European Commission to enhance the circularity of the plastics industry is ReBioCycle [139, 140], where robotic platforms are planned to be developed for waste sorting operations [140].
We began the design of circular flows of plastics by looking at several existing and potential configurations of the supply-recovery chain. Adopting the nomenclature introduced in Fig. 1 and the methodology outlined by Zocco et al. in [31], we conceptualize the five networks of thermodynamic compartments shown in Figs. 2(b)-2(f), where $\beta_{1}$ is the raw material to make the considered plastic (synthetic or bio-based) and $\beta_{2}$ is either synthetic plastic (networks in Figs. 2(b) and 2(c)) or bio-plastic (networks in Figs. 2(d), 2(e), and 2(f)). Since the compartmental networks result from a generalization of the Rankine cycle as explained in Section III, Fig. 2(a) shows the Rankine cycle, where $\beta_{1}$ is the working fluid. On the left-hand side, the cycle is depicted with the control surfaces indicated by dashed lines as in Section 8.2 of [121], while on the left-hand side, the cycle is conceived as a network of compartments, with each compartment indicated by its control surface (a rectangle with solid line). The control surfaces indicate the boundaries of each compartment and similarly in Figs. 2(b)-2(f). This way of representing sequences of interconnected compartments is called a “compartmental diagram” [31].
Figs. 2(b) and 2(c) consider a linear system and a more circular system for synthetic plastic, respectively. The red arrows indicate the unsustainable flows consisting of two types: raw materials extracted from non-renewable reservoirs entering the system and leaks of plastic leaving the system. The return flows into the supply chain are indicated with green arrows. As shown, both Fig. 2(b) and Fig. 2(c) have red flows in input and output, but in the former they are bigger because the latter retains a fraction of the plastic in the loop to meet the manufacturer demand $c^{2}_{2,2}$; the former has no return flows, while the latter has one return flow. The compartments that operate autonomously are depicted in yellow. The network in Fig. 2(c) sorts the waste in an autonomous facility ($c^{4}_{4,4}$), which produces two streams: the multi-component products are sent to an autonomous disassembly facility to recover the plastic parts, which are then sent to the plastic recycler; and in contrast, plastic items (e.g., bottles, packaging) are sent directly to the recycler. Typically, a fraction of the plastic is not recyclable due to contamination or quality degradation; this causes the leak of plastic to $c^{7}_{7,7}$.
The remaining three subfigures of Fig. 2 consider bio-plastic. Specifically, Fig. 2(d) shows an existing chain, which recovers bio-plastic through the recycling of bio-plastic products. The network has one red flow instead of two because of the consideration of bio-plastic instead of synthetic plastic. In Fig. 2(e), the bio-plastic components are disassembled from multi-component products and reused in products that differ from the initial ones due to the disassembly compartment ($c^{6}_{6,6}$), while in Fig. 2(f) bio-plastic components are reused in the original repaired products. Hence, while the former has two return flows to the manufacturer $c^{2}_{2,2}$ (the manufacturer uses the disassembled parts to make products), the latter has one return flow to the manufacturer and one to the retailer (the repaired products are ready to be sold again). Given that, to date, it is unlikely that multi-component products contain bio-plastic parts, Figs. 2(e) and 2(f) consider hypothetical compartmental configurations.
(a) Rankine cycle (left) conceived as a compartmental network (right)
(b) Network of synthetic plastic (linear architecture)
(c) Network of synthetic plastic (more circular architecture)
(d) Existing network of bio-plastic (no disassembly)
(e) Hypothetical net of bio-plastic (disassembly to reuse parts)
(f) Hypothetical net of bio-plastic (disassembly for repairing)
Figure 2: Initial application of the proposed framework to design plastic and bio-plastic flows. The framework is a generalization of the Rankine cycle (Fig. 2(a)), where the working fluid is designed to be circular. At this initial stage, we focus on the sorting and the disassembly compartments performed via robotic manipulators controlled with deep RL (compartments in yellow).
Each network in Fig. 2 can be represented by a set $\mathcal{N}$ of thermodynamic compartments [31]. For example, the system in Fig. 2(c) is represented by $\mathcal{N}_{1}$ as
$\mathcal{N}_{1}=\{c^{1}_{1,1},c^{2}_{2,2},c^{3}_{3,3},c^{4}_{4,4},c^{5}_{5,5},$ (1)
$c^{6}_{6,6},c^{7}_{7,7},c^{8}_{1,2},c^{9}_{2,3},c^{10}_{3,4},$
$c^{11}_{4,5},c^{12}_{4,6},c^{13}_{6,5},c^{14}_{5,7},c^{15}_{5,2}\}.$
Thus, the problem of designing circular flows can be stated as
$\mathcal{N}^{*}=\operatorname*{arg\,min}\,\,\,\dot{m}_{\text{u}}(\mathcal{N}),$ (2)
where $\dot{m}_{\text{u}}(\mathcal{N})$ is the sum of the unsustainable flows, i.e., the sum of the red flows, whose SI unit is kg/s. Equation (2) consists of finding the optimal network $\mathcal{N}^{*}$ that minimizes the unsustainable flows. Interventions on $\mathcal{N}$ to find $\mathcal{N}^{*}$ could be adding or removing compartments, modifying their interconnections, or changing their behaviour.
Remark.
Note the analogy between (2) and the optimization of an artificial neural network $\mathcal{N}_{\textup{n}}$, which can be stated as
$\mathcal{N}^{*}_{\textup{n}}=\operatorname*{arg\,min}\,\,\,L(\mathcal{N}_{\textup{n}}),$ (3)
where $L(\cdot)$ is a loss function [141]. There is, however, a key difference between $\mathcal{N}_{\textup{n}}$ and $\mathcal{N}$: the constituent unit of the former is, in its simplest form, a linear combination of the input and the weights filtered by an activation function, i.e., the neuron; in contrast, the constituent unit of the latter is, in its simplest form, a first-order ordinary differential equation derived from a mass balance or a thermodynamic law, i.e., the compartment.
In this initial study, we focus on the sorting and disassembly of compartments being carried out by robots. The simulation results for a two-link revolute-revolute (RR) planar manipulator controlled with different RL agents are reported in Table I. The tested algorithms are the deep deterministic policy gradient (DDPG) [142], the A2C [143], the proximal policy optimization (PPO) [144], and the soft actor-critic (SAC) [145]. For the DDPG algorithm, we used a multilayer perceptron as the policy model, a learning rate of 0.001, a batch size of 256, a soft update coefficient of 0.005, and a discount factor of 0.99. For the A2C algorithm, we used a multilayer perceptron as the policy model, a learning rate of 0.0007, a number of steps of 5, a discount factor of 0.99, an entropy coefficient of 0, a value function coefficient of 0.5, and the RMSprop optimizer. For the PPO algorithm, we used a multilayer perceptron as the policy model, a learning rate of 0.0003, a number of steps of 2048, a batch size of 64, a number of epochs of 10, a discount factor of 0.99, an entropy coefficient of 0, a value function coefficient of 0.5, and action noise exploration. Finally, for the SAC algorithm, we used a multilayer perceptron as the policy model, a learning rate of 0.0003, a batch size of 256, a soft update coefficient of 0.005, a discount factor of 0.99, a training frequency of 1, and an entropy regularization coefficient learned automatically.
The simulation was performed using the Python library MuJoCo [146]. The action of the robot environment is a vector whose elements are the joint torques, while there are 11 observation variables including the distance between the tip of the second link and the target point. The target point is generated at a random position at the beginning of every simulation.
TABLE I: Simulation performance of RL algorithms to control a two-link RR planar manipulator.
Algorithm Reward after training Training time (min:sec) Mean time per timestep (CPU; GPU) (sec) Success rate
A2C -10.99 4:56 0.00075; 0.0009 60.91%
DDPG -4.73 14:21 0.00058; 0.00071 93.37%
PPO -43.79 4:15 0.00073; 0.00092 41.83%
SAC -5.23 24:49 0.00079; 0.00102 95.97%
The success rate in Table I is calculated by considering a success whenever the tip of the second link reaches a distance from the target that is smaller than 4 cm and, simultaneously, the joint torques are smaller than 0.005 Nm (the reason for checking the torques along with the distance from the target is that the robot may reach the target and then, if the torques are high go beyond it, and thus failing the task; instead, the task is considered completed when the tip reaches the target and stops). The success rate is calculated over 10,000 simulations. The best algorithm is SAC followed by DDPG, although SAC requires also the longest time for training (24 min 49 sec). The next step will be to find a suitable RL algorithm to control a robot with at least 5 degrees of freedom performing pick-and-place in a waste sorting scenario, e.g., with a conveyor belt.
The simulations with the RR manipulator will also be used to analyse how the RL algorithm affects the whole network. For example, the robot controlled via the SAC agent could sort $q$ items per hour, which results in $m_{\text{r}}$ tonnes of plastic sorted in 1 day. In this case, $m_{\text{r}}$ tonnes of plastic are sent to the recycling center whose recycling time takes $t_{\text{r}}$ hours. If the manufacturer requires $m_{\text{r}}$ tonnes in less than $t_{\text{r}}$ hours, then the manufacturer may buy some virgin plastic or order from another recycler. Moreover, the failure rate of SAC (4.03%) indicates that some items will not be recycled resulting in a flow to the landfill that needs to be quantified and contained. The impact that the RL agent has on the supply-recovery chain can be captured only by looking at the whole network. This type of fundamental analysis regarding the material circularity is another important area of future work.
VI-B Limitations and Challenges
Since the framework is based on the unique universal science of thermodynamics [34], it is very general. This generality renders the framework useful to designing any type of material and, given the network formalization, highly scalable. However, this generality may yield a lack of focus and the rise of confusion about its purpose. By proposing the framework within a circular economy setting, we hope to provide the required clarity to the reader on the final goal that motivated its conceptualization. Another main challenge will be to create and maintain a two-way exchange of design ideas of the networks such as those in Fig. 2 between industry, with its practicality, and researchers, with their creativity. Without the former, the innovative ideas would remain in papers; without the latter, the possible pathways for a transition towards a circular society would be more difficult to envision.
On a more technical level, the first main challenge in implementing the framework to real scenarios is that the number of compartments to deal with could be large even for small supply-recovery chains because the materials and the product life-cycle stages in a modern society are many and highly dynamic. The second main challenge is to translate any network or compartment alterations into their impact on the unsustainable flows as compactly expressed by Equation (2). Specifically, if a compartment is, compartment-wise, far from any unsustainable flows, then the quantification of the impact of its dynamics on the unsustainable flows requires to back- and forward-propagate the variation to the following and proceeding compartments that separate it from the unsustainable flows. For example, the performance of the RL controller in the waste sorting facility ($c^{4}_{4,4}$) affects directly the neighboring compartments ($c^{12}_{4,6}$, $c^{11}_{4,5}$, and $c^{10}_{3,4}$). However, to quantify its impact on the red arrows it is necessary to propagate its impact on the intermediate compartments one at a time both forward ($c^{12}_{4,6}$, $c^{6}_{6,6}$, $c^{13}_{6,5}$, etc.) and backward ($c^{10}_{3,4}$, $c^{3}_{3,3}$, etc.) until the red flows are reached.
VII CONCLUSION